### v2

# Masked Language Models are Good Heterogeneous Graph Generalizers

Abstract
Heterogeneous graph neural networks (HGNNs) excel at capturing structural and semantic information in heterogeneous graphs (HGs), while struggling to generalize across domains and tasks. With the rapid advancement of large language models (LLMs), a recent study explored the integration of HGNNs with LLMs for generalizable heterogeneous graph learning. However, this approach typically encodes structural information as HG tokens using HGNNs, and disparities in embedding spaces between HGNNs and LLMs have been shown to bias the LLM’s comprehension of HGs. Moreover, since these HG tokens are often derived from node-level tasks, the model’s ability to generalize across tasks remains limited. To this end, we propose a simple yet effective Masked Language Modeling-based method, called MLM4HG. MLM4HG introduces metapath-based textual sequences instead of HG tokens to extract structural and semantic information inherent in HGs, and designs customized textual templates to unify different graph tasks into a coherent cloze-style <mask> token prediction paradigm. Specifically, MLM4HG first converts HGs from various domains to texts based on metapaths, and subsequently combines them with the unified task texts to form a HG-based corpus. Moreover, the corpus is fed into a pretrained LM for fine-tuning with a constrained target vocabulary, enabling the fine-tuned LM to generalize to unseen target HGs. Extensive cross-domain and multi-task experiments on four real-world datasets demonstrate the superior generalization performance of MLM4HG over state-of-the-art methods in both few-shot and zero-shot scenarios. Our code is available at https://github.com/BUPT-GAMMA/MLM4HG.
1 Introduction
Heterogeneous graphs (HGs) containing diverse node and edge types are ubiquitous in real-world scenarios, such as social networks Cai et al. (2024), knowledge graphs Chen et al. (2023), recommender systems Sun et al. (2019), and biological networks Ma et al. (2023). Embedding the structure and attribute information of HGs into a low-dimensional vector space, numerous heterogeneous graph neural networks (HGNNs) Fu et al. (2020); Wang et al. (2019b) and self-supervised learning (SSL) Tian et al. (2023); Wang et al. (2021) methods have been proposed and demonstrated promising performance in plenty of graph tasks. Nevertheless, these methods are primarily designed for a single HG or a specific graph task, significantly limiting their generalization across different domains and tasks Tang et al. (2024b). Consequently, the development of generalizable heterogeneous graph learning methods has emerged as a central focus Liu et al. (2023a).
With the great success of large language models (LLMs), researchers turn to harnessing LLMs for generalizable graph learning Tang et al. (2024a). To leverage the powerful understanding ability of LLMs and preserve the graph structural information, current researches either utilize GNNs to encode graphs into graph tokens Tang et al. (2024a) or converts graphs into texts Tan et al. (2024), after which the tokens or texts are fed into LLMs for inference. These approaches are mainly tailored to homogeneous graphs and overlook the rich structural and semantic information inherent in HGs. A recent research, HiGPT Tang et al. (2024b), introduces heterogeneity-aware graph instructions to help LLM deal with HGs. However, the structural information is extracted via an HGNN and encoded as HG tokens. The inherent mismatch between the embedding spaces of HGNNs and LLMs Lin et al. (2024) introduces biases into the LLM’s understanding of HGs. Moreover, HiGPT primarily focuses on node-level tasks, overlooking other essential graph tasks such as link prediction, which hinders its multi-task generalization ability and leads to suboptimal performance.
To address these problems, we propose a simple yet effective Masked Language Modeling-based method for generalizable Heterogeneous Graph learning, namely MLM4HG. To enhance cross-domain generalization, our method forgoes HGNNs and instead extracts the structural and semantic information of HGs from metapath-based sequences. Furthermore, to support multi-task generalization, we reformulate various graph tasks into a unified cloze-style <mask> prediction paradigm using customized textual templates. Specifically, we first employ a metapath-based graph-to-text method to extract the structural and semantic properties of HGs and convert them into texts. Next, we customize textual templates to reformulate different tasks into task sequences. These textual sequences are then combined to form an HG-based text corpus. Finally, after replacing the label tokens (e.g., ground truth node categories) with <mask> in the corpus, a pretrained masked language model is fine-tuned to predict the masked tokens using a constrained target vocabulary. The fine-tuned LM can subsequently be adapted to unseen target HGs across different graph tasks.
We conduct extensive cross-domain and multi-task experiments by fine-tuning the model on source HGs across different graph tasks and adapting it to an unseen target HG. The evaluation is performed on four real-world datasets from distinct domains and graph tasks in both few-shot and zero-shot scenarios. The results demonstrate that MLM4HG outperforms the best-performing baselines by an average of 18.08% in Micro-F1 and 7.10% in Macro-F1 in few-shot settings, as well as 15.22% in Micro-F1 and 8.04% in Macro-F1 in zero-shot settings. Our contributions are summarized as follows:
$\bullet$ To the best of our knowledge, we are the first to standardize HG representations and reframe different graph tasks as a unified cloze-style <mask> token prediction paradigm, establishing a generalizable learning framework fully powered by a masked language model.
$\bullet$ We propose MLM4HG, which first transforms HGs from various domains into texts based on metapaths, and subsequently combines them with the unified task texts to form a HG-based corpus. Then, the corpus is used to fine-tune a pretrained masked language model with a constrained target vocabulary to enable generalization.
$\bullet$ We conduct extensive cross-domain and multi-task experiments on four real-world datasets, demonstrating MLM4HG’s superior generalization performance across domains and tasks.
2 Related Work
2.1 Heterogeneous Graph Representation Learning
Heterogeneous graph neural networks (HGNNs) Wang et al. (2022); Yang et al. (2021) and self-supervised learning (SSL) Yang et al. (2022) are effective methods for heterogeneous graph representation learning.
Existing HGNNs can be broadly categorized into two types. Relation-based methods Hu et al. (2020b); Lu et al. (2019); Schlichtkrull et al. (2018) capture heterogeneous structural information by treating different edge types as distinct relations. For example, RHINE Lu et al. (2019) distinguish the various relations into Affiliation Relations (ARs) with one-centered by-another structures and Interaction Relations (IRs) with peer-to-peer structures. Considering the semantic information inherent in HGs, lots of metapath-based methods Dong et al. (2017); Fu et al. (2020); Wang et al. (2019b) have emerged in recent years. For instance, HAN Wang et al. (2019b) adopts hierarchical attention to aggregate information at both the node and semantic levels.
SSL methods include contrastive methods and generative methods. Contrastive methods Jiang et al. (2021); Wang et al. (2021) aim to maximize mutual information between different views. For example, HeCo Wang et al. (2021) employs network schema and meta-path views to assist the model in learning high-level node embeddings. Generative methods Hu et al. (2020a); Tian et al. (2023), on the other hand, leverage graph reconstruction and property prediction to extract general structural and attribute semantics. For instance, GPT-GNN Hu et al. (2020a) utilizes self-supervised attribute generation and edge generation tasks to capture the inherent dependencies between node attributes and graph structure. However, these methods focus on a single HG domain or a specific task, lacking the capacity to generalize across diverse HGs and tasks.
2.2 Large Language Models for Graphs
With the rapid advancements of large language models (LLMs) Brown et al. (2020); Touvron et al. (2023), researchers are increasingly focusing on LLM-enhanced graph learning Liu et al. (2023b). Recent approaches can be categorized into three main types. The first category utilizes GNNs to generate graph tokens, which aid LLMs in capturing structural information Chai et al. (2023); Liu et al. (2024); Tang et al. (2024a). The second category leverages LLMs to generate labels, features, and prompts for subsequent GNN training Li et al. (2024); Liu et al. (2023a); Ye et al. (2023). The third category attempts to use LLMs instead of GNNs as the backbone Chen et al. (2024); Lin et al. (2024); Zhu et al. , aiming to directly convert graphs into LLM-compatible sequences Tan et al. (2024) and utilize various instructions to assist LLMs in conducting downstream graph tasks. However, these approaches focus on homogeneous graphs and overlook the heterogeneous graphs prevalent in real-world scenarios. Although the recent HiGPT Tang et al. (2024b) makes an initial attempt at recognizing heterogeneity, its reliance on HGNNs introduces biases in the LLM’s understanding of HGs. Additionally, its focus on node-level tasks limits its ability to support multi-task learning, which further leads to suboptimal performance.
3 Preliminaries
This section first introduces basic concepts and then presents the learning scenarios to evaluate the generalization ability.
Definition 1. Heterogeneous Graph. A heterogeneous graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$. $\mathcal{V}$ represents the set of nodes with a node type mapping function $\varphi:\mathcal{V}\mapsto\mathcal{A}$, and $\mathcal{E}$ denotes the set of edges with an edge type mapping function $\psi:\mathcal{E}\mapsto\mathcal{R}$. $\mathcal{A}$ and $\mathcal{R}$ respectively denote the sets of node and edge types, where $|\mathcal{A}|+|\mathcal{R}|>2$. $\mathcal{X}$ represents a textual attribute set of nodes, and each node $v_{i}\in\mathcal{V}$ is associated with a textual attribute feature $x_{i}\in\mathcal{X}$.
Definition 2. Metapath. A metapath $P$ is defined as a path template in the form of $A_{1}\stackrel{{\scriptstyle R_{1}}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle R_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle R_{l}}}{{\longrightarrow}}A_{l+1}$, which describes a composite relation $R=R_{1}\circ R_{2}\circ\cdots\circ R_{l}$ between node types $A_{1}$ and $A_{l+1}$, where $\circ$ denotes the composition operator on relations.
Learning Scenarios. To evaluate the model’s generalization ability across various domains and tasks, following previous works Chen et al. (2024); Tang et al. (2024b), we conduct multi-task training on various source HGs and evaluate it on a target HG. The source and target HGs are from distinct domains. We refer to this configuration as the cross-domain and multi-task setting. In the target HG, only a few or even no nodes are labeled, which can be abstracted as few-shot or zero-shot scenarios. The few-shot scenario is often described as an N-way K-shot setting: N-way refers to the number of classes, and K-shot indicates that there are K labeled nodes per class, while in the zero-shot case, K is set to 0. Our goal is to extract generalizable knowledge under the cross-domain and multi-task setting, and subsequently transfer it to an unseen target HG across different graph tasks in both few-shot and zero-shot scenarios.
Figure 1: The overall framework of the proposed MLM4HG. We first convert HGs from different domains as well as the corresponding graph tasks into a unified textual format, forming a HG-based text corpus. Then we fine-tune a pretrained LM on the corpus by predicting masked label tokens. Finally, the LM can be adapted to unseen target HGs for generalization. Note that our performance does not solely rely on the prior knowledge lying in pretrained LMs. In Section 4.3, we anonymize nodes in the corpus by their ids, and can still achieve satisfying results.
4 Methodology
This section presents our proposed MLM4HG, with its framework shown in Figure 1. We first convert HGs from various domains into texts based on metapaths, and unify different graph tasks by textual templates. The above texts are then combined as a heterogeneous graph-based corpus. Afterward, we fine-tune a pretrained LM on the corpus by predicting the masked label tokens. Finally, the LM can be adapted to unseen target HGs for generalizable HG learning.
4.1 Heterogeneous Graph-based Corpus Construction
To help LMs generalize across different domains and tasks, we construct a heterogeneous graph-based corpus by converting relevant information into text sequences. Specifically, the HG-based corpus contains entries from different HGs and tasks, and each entry is composed of a task sequence and relevant nodes’ metapath-based sequences.
Metapath-based Sequences. Heterogeneous graphs contain rich structural and semantic information, which can be extracted through metapaths Fu et al. (2020); Wang et al. (2019b). For each node in a heterogeneous graph, metapath instances can be sampled based on its neighborhood, and subsequently textualized as metapath-based sequences. In this way, our method unifies the formats of different HGs, and enables the generalization ability across domains.
Specifically, we first initialize metapaths for each heterogeneous graph based on node and edge types. For each node $v$, we will consider all the metapaths starting with node type $\varphi(v)$, and then sample metapath instances in the form of $v\stackrel{{\scriptstyle e_{1}}}{{\longrightarrow}}v_{1}\stackrel{{\scriptstyle e_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle e_{l}}}{{\longrightarrow}}v_{l}$. Then each node $v_{i}$ in the metapath instance is textualized by concatenating its node type $\varphi(v_{i})$ and textual attribute $x_{i}$ as $T(v_{i})$, such as $<$ disease breasts $>$. Further, each metapath instance is textualized in the form $[T(v),\psi(e_{1}),T(v_{1}),...,T(v_{l-1}),\psi(e_{l}),T(v_{l})]$, where $\psi(e)$ is the edge type of $e$. Finally, all metapath instances associated with node $v$ are concatenated with special token </s> for separation, denoted as text sequence $MS_{v}$. For example, the metapath-based sequence for disease node breasts is illustrated in Figure 1.
Task Sequences. To enable the LM to generalize across different graph tasks, we unify task formats by the following textual templates:
In the classification task, the target is to predict the category of an unlabeled object. Therefore, we standardize the classification task format as $<$ You can deduce the category of [object] as <target>. $>$, where [object] refers to the textualized object to be classified and <target> represents the category label. For example, for the classification of disease node breasts, the corresponding task sequence is formatted as $<$ You can deduce the category of disease breasts as <glandular disease>. $>$.
In the link prediction task, the target shifts to determining whether an edge exists between two nodes and identifying its type. Therefore, we standardize the link prediction task format as $<$ You can deduce [src] <target> [dst]. $>$, where [src] and [dst] respectively represent the textualization $T(v_{1})$ and $T(v_{2})$ of source node $v_{1}$ and destination node $v_{2}$. <target> denotes the edge type between $v_{1}$ and $v_{2}$. For example, for the link prediction between the source node gene thyroid transcription factor-1 and the destination node disease bronchial abnormalities, the corresponding task sequence is formatted as $<$ You can deduce gene thyroid transcription factor-1 <causing> disease bronchial abnormalities. $>$. Utilizing these templates, we can automatically collect task sequences $TS$ covering all tasks and their examples.
Corpus Organization. Each entry in our HG-based corpus $\mathcal{I}$ consists of a specific task sequence with the metapath-based sequences of relevant nodes as prefix. Specifically, the $j$-th example of the $i$-th task is represented as entry $\mathcal{I}^{i}_{j}=[MS_{v_{1}},MS_{v_{2}},...,MS_{v_{M}},TS^{i}_{j}]$, where $M$ is the number of related nodes. For example, $M=1$ for the node classification task, and $M=2$ for link prediction.
4.2 Masked Language Modeling for Generalizable HG Learning
Given the HG-based corpus, we fine-tune a pretrained LM by multiple graph tasks via masked token prediction. Then the LM can generalize to tasks on unseen target HGs with limited annotations.
Masking and Tokenization. Different from the random masking strategy in BERT Devlin (2018), we only replace <target> tokens with <mask> for each entry. For example, in the node classification task for disease breasts, the target category <glandular disease> will be replaced with <mask>.
MLM with Constrained Vocabulary. The targets of graph tasks typically focus on a constrained set. For instance, in the node classification task of disease breasts shown in Figure 1, we aim to determine whether the <mask> token corresponds to a disease category, such as <cardiovascular disease>, <glandular disease>, or <cancer>, rather than considering the entire vocabulary as in typical language model tasks. Therefore, we adopt a constrained output vocabulary for each graph task. Specifically, each entry $\mathcal{I}^{i}_{j}$ is paired with a constrained target vocabulary $\mathcal{C}^{i}_{j}$, facilitating the LM to generalize across different datasets and tasks. This design choice closely mirrors the use of instruction-based output constraints in LLMs Tang et al. (2024a, b). For example, in the IMDB node classification task, the LLM is guided by the instruction: "Which of the following classes does this movie belong to: action, comedy, drama?"
We denote the masked token sequence of entry $\mathcal{I}^{i}_{j}$ as $\text{Tok}_{j}^{i}$, and the corresponding ground truth target as $\text{Tar}_{j}^{i}\in\mathcal{C}^{i}_{j}$. The optimization objective is defined by the classical cross entropy loss Devlin (2018); Sanh (2019) as:
$\mathcal{L}_{\text{MLM}}=-\frac{1}{|\mathcal{I}|}\sum_{i,j}\log p_{\text{MLM}}\left(\text{Tar}_{j}^{i}\mid\text{Tok}_{j}^{i}\right),$ (1)
where $p_{\text{MLM}}$ is the language modeling probability within the constrained vocabulary $\mathcal{C}^{i}_{j}$.
Generalization to Unseen Target HGs. To adapt the fine-tuned LM to unseen target HGs, we first transform the target HGs into the same format as the corpus. Then for zero-shot scenarios, we directly ask the LM to predict within the corresponding constrained vocabulary; for few-shot scenarios, we will update the parameters of LM in the same way as our fine-tuning process before inference. Detailed analyses of the scalability of MLM4HG are presented in Appendix D.
4.3 Discussion
Why Masked LMs Instead of LLMs? MLM4HG reformulates various tasks into a unified cloze-style <mask> token prediction paradigm, which essentially corresponds to classification. Prior work Zhang et al. (2024) shows that, compared to causal language modeling-based LLMs, masked LMs are better suited for such tasks both theoretically and empirically. Consequently, masked LMs align more naturally with the cloze-style formulation. Empirically, they achieve superior performance (Sections 5.2 and 5.3) while incurring significantly lower computational overhead (Appendix E).
Relying Solely on Pretrained LM’s Prior Knowledge? As discussed in this subsection, we adopt a pretrained language model with inherent prior knowledge as the backbone. To access whether the LM can effectively capture structural information embedded in metapath-based sequences, which is independent of its original knowledge, and generalize to the target dataset, we conduct an exploratory experiment on the supervised node classification task using the IMDB dataset.
As shown in Figure 2, we process the metapath-based sequences into three types: Attribute Only, Structure Only, and Both Attribute and Structure. In Attribute Only, only attribute information is included. In Structure Only, all nodes are anonymized with no attribute information, while in Both Attribute and Structure, both structure and attribute information are provided. The results are shown in Figure 2. In the 20% supervised learning scenarios, all three input types outperform the metapath-based HGNN baseline HAN Wang et al. (2019b), with Both Attribute and Structure achieving the best results. This highlights the effectiveness of metapath-based sequences in allowing the LM to integrate both structural and semantic information from various source heterogeneous graphs. While the LM’s prior knowledge contributes to the strong performance of Attribute Only, the competitive results of Structure Only indicate that our approach does not merely rely on pretrained knowledge. Instead, the LM is able to extract and leverage structural signals from metapath-based sequences for downstream tasks.
Figure 2: Metapath-based sentences in (a) Attribute Only, (b) Structure Only and (c) Both Attribute and Structure settings. Experiments are conducted on the IMDB node classification task.
5 Experiments and Analysis
To validate the effectiveness of our MLM4HG model, we conduct extensive cross-domain and multi-task experiments on four real-world heterogeneous graph datasets from various domains under diverse settings to answer the following research questions:
RQ1: How well does MLM4HG generalize across domains in few-shot scenarios?
RQ2: How well does MLM4HG generalize across tasks in zero-shot scenarios?
RQ3: What is the impact of model components and learning scenarios on performance?
RQ4: How effective are metapath-based sequences compared to other graph-to-text methods?
Additionally, we further investigate the impact of using different LMs or LLMs as backbones in Appendix E. We also analyze the impact of different hyperparameters (Appendix F) and provide a case study (Appendix I) with a visual analysis of attention scores from the final layer of MLM4HG across various graph tasks.
5.1 Experimental Settings
Datasets. We conduct extensive experiments on four real-world heterogeneous graph benchmark datasets from different domains: IMDB, DBLP Wang et al. (2019b), YELP, and PubMed Yang et al. (2020). More details can be found in Appendix A.
Baselines. To comprehensively evaluate the performance of our MLM4HG, we compare it with the following five representative and state-of-the-art methods from three different categories. The first category is represented by the HGNN method HAN Wang et al. (2019b). The second category is the heterogeneous graph self-supervised learning method HeCo Wang et al. (2021). The third category includes LM and LLM-enhanced graph methods, including WalkLM Tan et al. (2024), GraphGPT Tang et al. (2024a), and HiGPT Tang et al. (2024b), in which HiGPT is specifically designed for heterogeneous graphs, while the remaining methods are tailored for homogeneous graphs.
To prevent potential concerns regarding fairness due to our use of constrained vocabularies, we provide additional clarification here. HiGPT and GraphGPT restrict the output space of LLMs through instructions such as: "Which class does this movie belong to: action, comedy, drama?", a strategy that successfully bounds over 99% of the generated outputs. Similarly, other baselines adopt task-specific prediction heads that inherently limit the output dimensions. These practices collectively ensure a fair basis for comparison. Further details are provided in Appendix B.
Evaluation Protocols. We adopt the cross-domain and multi-task setting during fine-tuning. For instance, when IMDB serves as the target HG, we fine-tune the pretrained language model on DBLP, PubMed, and YELP as source HGs, using both node classification and link prediction tasks. The model’s generalization ability is then evaluated on the unseen target HG across these two tasks. For the node classification task, we evaluate on both few-shot and zero-shot scenarios. The few-shot setting follows an N-way K-shot protocol with K in {1, 5, 20, 40} Tang et al. (2024b). For the link prediction task, we assess model performance under the zero-shot scenario. Since we unify different graph tasks into the <mask> token prediction paradigm, the link prediction task in our experiments is essentially treated as an edge classification task. Accordingly, both node classification and link prediction are evaluated using Micro-F1 and Macro-F1 scores. (We also report AUC and AP metrics for the link prediction task in Appendix G.) All reported results are averaged over five independent runs.
Implementation Details. The LM used in our experiments is pretrained DistilRoBERTa Sanh (2019). To automatically generate diverse metapaths while reducing reliance on manual design, we comprehensively traverse all edge types to construct candidate metapaths, as shown in Appendix A. For each node in HGs, we randomly sample 3 instances per metapath. This design choice is motivated by our observation (Appendix F) that longer metapath-based sequences enable the LM to better capture intricate structural and semantic information. Notably, sampling three instances yields sequence lengths that closely align with the model’s maximum input capacity of 512 tokens, thereby fully utilizing its contextual modeling capabilities. The LM is fine-tuned for three epochs, and the other hyperparameters of MLM4HG are optimized through grid-searching for best performance. The baseline parameters are initially set according to the values reported in the original papers and then optimized through grid-searching to achieve optimal performance. All hyperparameter settings are provided in Appendix C, and a detailed hyperparameter study is available in Appendix F.
5.2 Cross-domain Generalization Analysis (RQ1)
To validate the model’s generalization ability on datasets from distinct domains, we fine-tune the model under the cross-domain and multi-task setting and then perform node classification experiments on target HGs in few-shot scenarios.
The overall experimental results are presented in Table 1. (1) MLM4HG consistently outperforms a range of state-of-the-art baselines across 16 evaluation groups on four datasets, achieving an average gain of 18.08% in Micro-F1 and 7.10% in Macro-F1. (2) Notably, MLM4HG with only 1-shot supervision surpasses other methods with 40-shot settings on the IMDB and PubMed datasets, indicating its strong ability to capture transferable knowledge from source HGs and effectively generalize to target HGs with minimal supervision. (3) HiGPT achieves the second-best performance in most cases and outperforms GraphGPT, benefitting from its heterogeneous graph instruction tuning, which helps it recognize relation types more effectively. However, the reliance on HG tokens generated by HGNNs introduces embedding space mismatch, and the lack of metapath-level semantics further hinders its performance. (4) WalkLM performs better than HiGPT on PubMed, likely because its random walk-based sequences enable the LM to more directly capture graph structures, which is especially useful for PubMed’s complex biomedical data. However, WalkLM fails to account for generalization and omits metapaths, which ultimately leads to suboptimal performance.
5.3 Multi-task Generalization Analysis (RQ2)
To further examine the model’s ability to generalize to different tasks on unseen target HGs, we perform node classification and link prediction tasks across four different datasets in zero-shot scenarios. We also fine-tune the model under the cross-domain and multi-task setting, followed by evaluation of node classification and link prediction tasks on the unseen target HG.
The results are summarized in Table 2. (1) Our MLM4HG outperforms baselines across the majority of datasets and tasks, demonstrating strong generalization capabilities across different graph tasks on unseen target HGs. While it initially exhibits suboptimal and random performance on the node classification task in DBLP, MLM4HG rapidly adapts to the domain of the target HG after 1-shot fine-tuning in Table 1, leading to notable improvements in performance. (2) In comparison to the baselines, MLM4HG achieves a significant enhancement in link prediction, with average gains of 27.81% in Micro-F1 and 13.57% in Macro-F1. This highlights the effectiveness of metapath-based sequences in enhancing the LM’s ability to capture intricate structural information of HGs, leading to significant improvements in link prediction performance.
Table 1: Average Micro-F1 and Macro-F1 scores for node classification in few-shot scenarios. The best results are in bold and the second best are underlined. Our results are significantly better than the best baseline, as confirmed by the Student’s t-test at the 0.01 significance level.
Metrix Micro-F1 Macro-F1
Target k-shot HAN HeCo WalkLM GraphGPT HiGPT MLM4HG HAN HeCo WalkLM GraphGPT HiGPT MLM4HG
IMDB 1-shot 0.3521 0.3401 0.3871 0.3671 0.4901 0.6529 0.3126 0.3027 0.3507 0.2782 0.3621 0.3973
5-shot 0.3622 0.3423 0.3701 0.3996 0.5849 0.6690 0.3117 0.2977 0.3448 0.3305 0.3817 0.4008
20-shot 0.3824 0.3681 0.4248 0.4061 0.5809 0.7032 0.3356 0.3098 0.3676 0.3461 0.3749 0.4129
40-shot 0.4002 0.3907 0.4337 0.4186 0.6076 0.7201 0.3527 0.3217 0.4121 0.3591 0.3976 0.4186
DBLP 1-shot 0.2538 0.2401 0.2542 0.2831 0.2622 0.2870 0.2435 0.2398 0.2452 0.2497 0.2437 0.2530
5-shot 0.2622 0.2537 0.2634 0.3152 0.3743 0.3925 0.2462 0.2404 0.2574 0.2563 0.2754 0.2819
20-shot 0.2839 0.2673 0.3093 0.3277 0.3674 0.6350 0.2680 0.2551 0.2988 0.3171 0.3379 0.3884
40-shot 0.2903 0.2894 0.3176 0.3301 0.3824 0.6817 0.2742 0.2635 0.2799 0.3205 0.3577 0.4054
PubMed 1-shot 0.1527 0.1624 0.1763 0.1916 0.1932 0.5661 0.1224 0.1312 0.1313 0.1323 0.1268 0.3615
5-shot 0.2576 0.2793 0.3495 0.2231 0.2659 0.6203 0.1946 0.2051 0.2284 0.1737 0.1743 0.3828
20-shot 0.2603 0.2832 0.3593 0.2265 0.2754 0.6508 0.2275 0.2307 0.2382 0.1731 0.1696 0.3943
40-shot 0.2834 0.2931 0.3659 0.2438 0.3037 0.6712 0.2423 0.2496 0.2662 0.1933 0.1807 0.4016
YELP 1-shot 0.1427 0.1135 0.1613 0.3164 0.4133 0.4353 0.0981 0.0879 0.1242 0.1856 0.2833 0.3033
5-shot 0.1832 0.1427 0.2145 0.3628 0.5413 0.6755 0.1513 0.1203 0.2069 0.2271 0.3431 0.4032
20-shot 0.2746 0.2294 0.2915 0.3836 0.5452 0.7524 0.2109 0.1632 0.2314 0.2336 0.3486 0.4294
40-shot 0.2831 0.2475 0.3023 0.3906 0.5713 0.7892 0.2379 0.2070 0.2401 0.2765 0.3624 0.4411
Table 2: Average Micro-F1 and Macro-F1 scores for node classification (NC) and link prediction (LP) in zero-shot scenarios. The best results are in bold and the second best are underlined.
Metric Micro-F1 Macro-F1
Task Target HAN HeCo WalkLM GraphGPT HiGPT MLM4HG HAN HeCo WalkLM GraphGPT HiGPT MLM4HG
NC IMDB 0.3233 0.3203 0.3456 0.3525 0.3895 0.4622 0.2201 0.2275 0.2352 0.2411 0.2838 0.3161
DBLP 0.2405 0.2437 0.2483 0.2465 0.2371 0.2477 0.1594 0.1600 0.1638 0.1659 0.1532 0.1601
PubMed 0.1435 0.1324 0.1536 0.1601 0.1811 0.1898 0.0827 0.0601 0.0514 0.1032 0.1232 0.1595
YELP 0.0635 0.0721 0.0798 0.1138 0.1532 0.1772 0.0273 0.0298 0.0399 0.0772 0.1134 0.1505
LP IMDB 0.1625 0.1542 0.1930 0.2066 0.3662 0.9054 0.0814 0.0793 0.0816 0.1176 0.1868 0.4752
DBLP 0.1236 0.1149 0.1397 0.2923 0.3246 0.5356 0.0779 0.0683 0.0843 0.2403 0.2355 0.3488
PubMed 0.1764 0.1527 0.1830 0.1866 0.1912 0.1986 0.0825 0.0673 0.0948 0.1125 0.1352 0.1657
YELP 0.1301 0.1232 0.1456 0.2491 0.4829 0.8378 0.0649 0.0587 0.0766 0.0881 0.3404 0.4559
5.4 Ablation Study (RQ3)
In this subsection, we introduce four variants to examine the influence of different learning scenarios and model components on the experimental results. To capture more comprehensive and general graph knowledge from source HGs, MLM4HG adopts cross-domain and multi-task fine-tuning by default. To evaluate whether the model benefits from this learning scenarios, we define two variants: w/o Fine-tuning and w/o Multi-task. w/o Fine-tuning indicates that the model is directly applied to the target HG without cross-domain and multi-task fine-tuning on source HGs. w/o Multi-task refers to fine-tuning the pretrained LM on source HGs using the single same task as the target HG. That is, w/o Multi-task adds multi-domain pretraining to w/o Fine-tuning, whereas the complete MLM4HG includes both multi-domain and multi-task pretraining. Additionally, to assess the impact of different components on the experimental outcomes, we design two variants: w/o Metapath and w/o Constraint. In w/o Metapath, the metapath-based sequences is replaced with node attributes. In w/o Constraint, the predicted <mask> token is selected from the full vocabulary rather than the constrained target vocabulary.
We evaluated the performance of these variants in node classification and link prediction tasks in both few-shot and zero-shot scenarios. The experimental results are shown in Table 3. (1) It is evident that these variants perform suboptimally compared to the original MLM4HG across different graph tasks and experimental settings. (2) w/o Multi-task outperforms w/o Fine-tuning and achieves competitive results overall, demonstrating that multi-domain pretraining enables the LM to extract transferable knowledge and significantly boosts performance, even when fine-tuned on a single task. Further performance gains achieved by MLM4HG highlight the benefit of multi-task fine-tuning, which facilitates the extraction of more comprehensive information from multiple perspectives. (3) When the vocabulary size is unconstrained in w/o Constraint, the results experience a significant decline compared to MLM4HG. This clearly indicates that the model’s focus should be directed toward the target categories, especially for our simpler graph tasks. (4) When node attributes replace metapath-based sequences in w/o Metapath, the performance of both node classification and link prediction tasks also decreases to varying degrees. This confirms that metapath-based sequences contains rich structural and semantic information, which helps the LM extract general heterogeneous graph knowledge and generalizing across different HGs. (5) We further observe that in the few-shot scenarios, all variants—except for w/o Constraint—perform comparably to MLM4HG. This is because the pretrained LM can quickly adapt to a small number of labeled examples, as shown in Table 1. For instance, our methods already outperform the strongest baselines under the 5-shot setting on the IMDB and YELP datasets. However, in the more challenging zero-shot scenario, each component becomes essential.
Table 3: Average Micro-F1 and Macro-F1 scores for MLM4HG variants on node classification and link prediction tasks over IMDB and YELP datasets.
Task Node Classification Link Prediction
k-shot 5-shot zero-shot zero-shot
Metric Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1
Target IMDB YELP IMDB YELP IMDB YELP IMDB YELP IMDB YELP IMDB YELP
w/o Fine-tuning 0.6497 0.6604 0.3875 0.3977 0.3666 0.0746 0.2682 0.0695 0.5254 0.1781 0.3444 0.1511
w/o Multi-task 0.6570 0.6701 0.3965 0.4012 0.4598 0.1476 0.3150 0.1286 0.8853 0.8162 0.4696 0.4494
w/o Metapath 0.6380 0.4917 0.3795 0.3296 0.4304 0.1519 0.2867 0.1389 0.8705 0.3173 0.4654 0.2409
w/o Constraint 0.3643 0.0547 0.2670 0.0519 0.0491 0.0000 0.0468 0.0000 0.4302 0.1772 0.3008 0.1505
MLM4HG 0.6690 0.6755 0.4008 0.4032 0.4622 0.1772 0.3161 0.1505 0.9054 0.8378 0.4752 0.4559
5.5 Graph-to-Text Method Comparison (RQ4)
We compare our metapath-based sequences with commonly used graph-to-text methods, specifically GML Himsolt (1997) and GraphML Brandes et al. (2013). Networkx provides a unified parser for these two formats, thus we use the parsed node and edge information as a replacement for the metapath-based sequences. Examples and parsing results of these two formats are provided in Appendix J.
As shown in Table 4, metapath-based sequences demonstrate superior performance in both node classification and link prediction tasks on the IMDB and YELP datasets. Compared to metapath-based sequences, GML and GraphML merely offer a straightforward description of nodes and edges, requiring the LM to further extract the rich and implicit semantic information, which increases the task complexity. This confirms that our metapath-based sequences effectively extracts both structural and semantic information from HGs, benefiting downstream tasks.
Table 4: Average Micro-F1 and Macro-F1 scores for node classification and link prediction using different graph-to-text methods on IMDB and YELP under both few-shot and zero-shot scenarios.
Task Node Classification Link Prediction
k-shot 5-shot zero-shot zero-shot
Metric Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1
Target IMDB YELP IMDB YELP IMDB YELP IMDB YELP IMDB YELP IMDB YELP
GML & GraphML 0.6337 0.5764 0.3879 0.3656 0.4043 0.0701 0.2879 0.0665 0.8984 0.6406 0.4587 0.3059
Metapath-based sequences 0.6690 0.6755 0.4008 0.4032 0.4622 0.1772 0.3161 0.1505 0.9054 0.8378 0.4762 0.4559
6 Conclusion
In this paper, we propose a simple yet effective Masked Language Modeling-based approach for generalizable Heterogeneous Graph learning, called MLM4HG. MLM4HG introduces metapath-based sequences to unify the representation of HGs, enabling cross-domain generalization. Then it reformulates different tasks into a cloze-style <mask> token prediction paradigm using customized templates, thereby supporting multi-task generalization. By constructing an HG-based corpus and using it to fine-tune a pretrained LM, we enable the LM to generalize across unseen target graphs and tasks. To the best of our knowledge, we are the first to apply masked language models for generalizable heterogeneous graph learning, establishing a new paradigm and paving a new path for future research. Extensive experiments in four real-world datasets from distinct domains demonstrate MLM4HG’s superior performance across different graph tasks in both few-shot and zero-shot scenarios. Limitations and broader impacts are discussed in Appendix H.
Appendix A Dataset Statistics
Table 5 summarizes dataset statistics, including node/edge types (with the target node type in bold), target categories, and metapaths used in the metapath-based sequence construction.
Table 5: Statistics of datasets.
Dataset Node Type Edge Type Target Node Category Metapath
IMDB
Movie (M), 4,278
Actor (A), 5,257
Director (D), 2,081
acted in, 12,828
was acted by, 12,828
directed, 4,278
was directed by, 4,278
Action
Comedy
Drama
MAM, MDM
AMA, AMDMA
DMD, DMAMD
DBLP
Author (A), 4,057
Paper (P), 14,328
Term (T), 7,723
Conference (C), 20
write, 19,645
was written by, 19,645
publish, 85,810
was published in, 85,810
receive, 14,328
was received by, 14,328
1, 2, 3, 4
APA, APCPA, APTPA
PAP, PTP, PCP
TPT, TPCPT, TPAPT
CPC, CPAPC, CPTPC
YELP
Business (B), 7,474
Location (L), 39
Phrase (P), 74,943
Stars (S), 9
rate, 7,474
score, 7,474
indicate, 2,654,313
located in, 7,474
described with, 2,654,313
is associated with, 7,474
beauty & spas
pizza, food
sandwiches
nightlife
burgers
Mexican
American (new)
shopping
automotive
Italian, bars
restaurants
breakfast & brunch
American (traditional)
event planning & services
BLB, BSB, BLB
LBL, LBPBL, LBSBL
SBS, SBLBS, SBPBS
PBP, PBLBP, PBSBP
PubMed
Disease (D), 20,163
Gene (G), 13,562
Chemical (C), 26,522
Species (S), 2,863
in, 97,297
with, 8,400
contain, 88,897
causing, 25,962
is caused by, 25,962
cardiovascular disease
glandular disease
nervous disorder
communicable disease
inflammatory disease
pycnosis
skin disease
cancer
DGD, DCD, DSD
GDG, GCG, GSG
CDC, CGC, CSC
SDS, SGS, SCS
Appendix B Baseline Descriptions
Detailed descriptions of the baselines are as follows:
  • HAN Wang et al. (2019b) develops a hierarchical attention structure that includes node-level attention and semantic-level attention to fully consider the importance of node and metapath. However, the end-to-end training manner limits its ability to support generalizable heterogeneous graph learning.
  • HeCo Wang et al. (2021) employs cross-view contrastive mechanism, where the two views of heterogeneous graph, including network schema and metapath views, collaboratively supervise each other and finally learn high-level node embeddings. Although it performs well in few-shot scenarios, its reliance on pretraining and fine-tuning for a specific heterogeneous graph significantly restricts its ability to generalize across different domains and tasks.
  • WalkLM Tan et al. (2024) proposes leveraging random walk-based textual sequences to capture both complex attribute information and flexible topological structures of graphs in an unsupervised manner. However, it overlooks the rich semantic information in HGs.
  • GraphGPT Tang et al. (2024a) proposes a dual-stage instruction tuning paradigm that aligns LLMs with graph structural knowledge, enabling generalization across unseen graphs in node classification tasks. However, GraphGPT only considers homogeneous graphs, neglecting the rich structural and semantic information present in HGs.
  • HiGPT Tang et al. (2024b) is a heterogeneous graph instruction-tuning model, which introduces heterogeneity-aware graph instructions to help the LLM comprehend the complex relations in HGs. However, HiGPT leverages HGNNs to extract structural information as HG tokens, leading to biases in LLMs’ understanding. Moreover, the node-level pretraining task limits its multi-task generalization ability, which further leads to suboptimal performance.
Appendix C Implementation Details
Our experiments are carried out on a Linux OS with a single NVIDIA A800 GPU (80GB memory). The MLM4HG implementation relies on libraries including PyTorch Paszke et al. (2019), PyTorch Geometric Fey & Lenssen (2019), DGL Wang et al. (2019a), scikit-learn Pedregosa et al. (2011), and Transformers Wolf et al. (2020). The model is fine-tuned for 3 epochs with a batch size of 16. During fine-tuning, the maximum token sequence length for input data is set to 512. The optimizer used is AdamW Loshchilov (2017), with a learning rate of 1e-5 and a weight decay of 0.01.
Appendix D Scalability Analysis
We discuss the time complexity and scalability of MLM4HG. MLM4HG employs a masked language model as the backbone, with time complexity scaling linearly with the token count in the corpus. For larger-scale graphs, longer metapath-based sequences may increase token sequence length and impact computational efficiency. To mitigate this, we randomly sampled 3 instances per metapath per node for the YELP dataset, which contains millions of edges, while still achieving promising results (Tables 1 and 2). Additionally, we provide detailed statistics in Table 6, including the number of nodes and edges, total token count (for both NC and LP), fine-tuning time per graph and inference time per batch (40 examples) across four datasets, further demonstrating MLM4HG’s scalability.
Table 6: Scalability analysis.
Node Num. Edge Num. Fine-Tuning Inference NC Token Count LP Token Count
IMDB 11,616 34,212 338.68s 1.0521s 934,956 1,810,104
DBLP 26,128 239,566 381.69s 1.3190s 1,684,221 2,458,781
PubMed 63,110 246,518 452.92s 1.2754s 79,015 1,291,117
YELP 82,465 5,338,522 334.44s 1.5963s 1,364,293 1,794,996
Appendix E Impact of Different LMs and LLMs
The impact of different LMs is worth investigating. Considering the recent advances of causal language modeling (CLM) approachesAchiam et al. (2023); Touvron et al. (2023), we not only evaluate MLM-based models, such as RoBERTa Liu (2019) and DistilRoBERTa Sanh (2019), but also explore the performance of LLaMA3 Touvron et al. (2023) (specifically, LLaMA-3.1-8B-Instruct) and Qwen3 Yang et al. (2025) (specifically, Qwen3-8B) in our experiments. To ensure a fair comparison, we leverage instructions to restrict CLM-based models’ predictions to the constrained target vocabulary. We also report the training time per batch for each model, with a batch size of 16. For MLM-based language models, full parameter fine-tuning are applied, while CLM-based models are fine-tuned using LoRA Hu et al. (2021).
Table 7: Average Micro-F1 scores for node classification and link prediction on IMDB and YELP with different LMs and LLMs. The best results are highlighted in bold and the second-best are underlined.
Task Node Classification Link Prediction
LM Time IMDB YELP IMDB YELP
DistilRoBERTa 0.47s 0.4622 0.1772 0.9054 0.8378
RoBERTa 0.68s 0.4417 0.2268 0.8785 0.7713
LLaMA3 2.28s 0.3364 0.1693 0.8584 0.6903
Qwen3 2.02s 0.2644 0.1475 0.6339 0.6013
We evaluated the impact of different LMs on the experimental results in node classification and link prediction tasks on the IMDB and YELP datasets. The experimental results are shown in Table 7. Language models trained with MLM outperform those using CLM and require less training time. While CLM excels in many generative tasks Wang et al. (2024), MLM is better suited for our classical graph tasks, which are unified into a cloze-style format and thus demand a stronger focus on both left-to-right and right-to-left contextual information. In addition, as a lightweight variant of RoBERTa, DistilRoBERTa achieves competitive results while significantly reducing both the parameter count and training time. Therefore, we adopt DistilRoBERTa as the backbone model in our experiments.
Appendix F Hyperparameter Study
We conducted a detailed evaluation of two key hyperparameters: maximum token sequence length (Max Length) and the number of fine-tuning epochs (Fine-tuning epochs), to understand their influence on the performance of MLM4HG. As illustrated in Figure 3, extending the token sequence length consistently improves results for both node classification and link prediction tasks. This suggests that longer metapath-based sequences enable the language model to capture richer structural and semantic context from the heterogeneous graph. Furthermore, increasing the number of fine-tuning epochs initially leads to noticeable performance gains for both tasks, after which the results tend to stabilize. Based on this observation, we fixed the number of fine-tuning epochs to 3 in all reported experiments to balance performance and training efficiency.
Figure 3: Average Micro-F1 and Macro-F1 scores for node classification (NC) and link prediction (LP) with different hyperparameter settings on IMDB dataset in the zero-shot scenarios.
Appendix G Zero-shot Link Prediction Evaluation Using AUC and AP
Table 8 presents the evaluation results of AUC and AP metrics, which serve as complementary measures for the zero-shot link prediction task. These results consistently exhibit the same trends as the F1 scores in Table 2, further validating our conclusions.
Table 8: Average AUC and AP scores for link prediction in zero-shot scenarios. The best results are highlighted in bold and the second-best are underlined.
Metric AUC AP
Target HAN HeCo WalkLM GraphGPT HiGPT MLM4HG HAN HeCo WalkLM GraphGPT HiGPT MLM4HG
IMDB 0.5194 0.5047 0.5438 0.5836 0.6416 0.9471 0.1892 0.1724 0.2067 0.2236 0.2822 0.7812
DBLP 0.5423 0.5302 0.5634 0.5942 0.6208 0.8160 0.1475 0.1325 0.1555 0.2292 0.2255 0.4762
PubMed 0.5167 0.5106 0.5242 0.5366 0.5452 0.6886 0.1506 0.1397 0.1657 0.1997 0.1858 0.3023
YELP 0.5143 0.5079 0.5237 0.5924 0.6967 0.8596 0.1296 0.1236 0.1451 0.2084 0.3630 0.6254
Appendix H Limitations and Broader Impacts
We discuss the limitations and broader impacts of MLM4HG in this section. Although our exploratory experiments on metapath-based sequences in Section 4.3 demonstrate that MLM4HG is able to extract and leverage structural signals from such sequences for downstream tasks, the semantic information encoded in textual content remains crucial for language models. Therefore, applying LMs to non-textual graph data still requires further exploration. Moreover, additional investigation is needed to assess the model’s generalizability across a wider range of graph tasks.
MLM4HG is a simple yet effective framework for generalizable heterogeneous graph learning. Its potential impact extends to a variety of real-world applications involving HGs, such as social network analysis, knowledge representation, and recommender systems. While the ability to generalize across domains and tasks can enhance the development of robust AI systems, it may also introduce concerns. For example, the model could be misused in illegal activities, leading to unintended consequences such as the perpetuation of biases or the misuse of predictive insights. We encourage future work to focus on enhancing the transparency and fairness of graph language models, especially in sensitive applications.
Appendix I Case Study
In this section, we first provide some entries of the heterogeneous graph-based corpus used in our experiments. We then perform a visual analysis of the attention scores computed by the last layer of the model. Higher scores correspond to darker background colors, indicating that the relevant words have a greater impact on the prediction of <mask>.
The entries for node classification and link prediction tasks are presented in Table 9. Fine-tuning the masked LM on such entries enables generalization across domains and graph tasks. Furthermore, we present a visualization of the attention scores computed by MLM4HG on the IMDB dataset. (1) Figure 4 (a) shows the visualization results of node classification task for the movie I Spy. From the intensity of the colors, we observe that the model first focuses on the node’s attribute information (name, I Spy) and node type (movie). Additionally, the model attends to information from neighboring nodes (28 Days and Doctor Dolittle), which belong to the same category as the target node. Finally, the model identifies the task as a node classification task (the category of), assisting in making the correct prediction. (2) Figure 4 (b) illustrates the link prediction task between the source node business The Prime Chinese Restaurant and the destination node phrase dog meat. In addition to the attributes of the source and destination nodes, the model places more emphasis on the structural information within the metapath sequences, such as <business described with phrase>, <phrase indicates business>, and <location is associated with business>, ultimately providing an accurate prediction.
Figure 4: Visualization of the attention scores computed by the last layer of our MLM4HG for different graph tasks. Higher scores correspond to darker background colors, indicating that the relevant words have a greater impact on the prediction of the <mask> token.
Table 9: Entries of the heterogeneous graph corpus for node classification (NC) and link prediction (LP) tasks.
Task Dataset Entries
NC IMDB For movie I Spy: [movie I Spy was acted by actor Phill Lewis acted in movie I Spy </s> movie I Spy was acted by actor Gary Cole acted in movie One Hour Photo </s> movie I Spy was directed by director Betty Thomas directed movie 28 Days </s> movie I Spy was directed by director Betty Thomas directed movie Doctor Dolittle </s>] You can deduce the category of movie I Spy as <comedy>.
DBLP For author Keith L. Doty: [author Keith L. Doty write paper MICRONET: A Microcomputer Network System for Managing Distributed Relational Databases. was published in term managing publish paper Managing Healthcare Data Hippocratically. was written by author Ameet Kini </s> author Keith L. Doty write paper Extending Object-Oriented Concepts to Support Engineering Applications. was written by author Keith L. Doty </s>] You can deduce the category of author Keith L. Doty as <0>.
PubMed For disease dermatoses: [disease dermatoses contain chemical trendline in disease tyrosination </s> disease dermatoses contain chemical F16357 in disease MDCP </s> disease dermatoses is caused by gene MicroRNA-146a causing disease dermatoses </s> disease dermatoses is caused by gene DmPpt1 causing disease autoimmune cytopenias </s>] You can deduce the category of disease dermatoses as <skin disease>.
YELP For business The Living Room: [business The Living Room described with phrase 15k miles indicate business Hollywood Beauty Eyebrow Threading </s> business The Living Room described with phrase mexican street corn indicate business Burger Theory </s> business The Living Room located in location Scottsdale is associated with business Daphne’s California Greek </s> business The Living Room rate stars 4.0 score business Bell Lexus North Scottsdale </s>] You can deduce the category of business The Living Room as <sandwiches>.
LP IMDB For movie Pixels: [movie Pixels was acted by actor Josh Gad acted in movie The Angry Birds Movie </s> movie Pixels was acted by actor Josh Gad acted in movie Ice Age: Continental Drift </s> movie Pixels was directed by director Chris Columbus directed movie Bicentennial Man </s>] For actor Adam Sandler: [actor Adam Sandler acted in movie Grown Ups was directed by director Dennis Dugan directed movie You Don’t Mess with the Zohan was acted by actor Kevin Nealon </s> actor Adam Sandler acted in movie Hotel Transylvania was acted by actor Steve Buscemi </s>] You can deduce movie Pixels <was acted by> actor Adam Sandler.
DBLP For term relational: [term relational publish paper Integrating Pattern Mining in Relational Databases. was written by author Toon Calders write paper Mining rank-correlated sets of numerical attributes. was published in term correlated </s> term relational publish paper Updating Relational Databases through Object-Based Views. was published in term updating </s>] For paper Learning Probabilistic Relational Models.: [paper Learning Probabilistic Relational Models. was published in term relational publish paper Using relational knowledge discovery to prevent securities fraud. </s> paper Learning Probabilistic Relational Models. was received by conf IJCAI receive paper Parametric Kernels for Sequence Data Analysis. </s>] You can deduce term relational <publish> paper Learning Probabilistic Relational Models. .
PubMed For gene U2af1-rs1: [gene U2af1-rs1 </s>] For disease neuroendocrine tumours: [disease neuroendocrine tumours contain chemical ammonium phosphate in disease neuroendocrine tumours </s> disease neuroendocrine tumours contain chemical DiMeIQx in disease chronic hypertension </s> disease neuroendocrine tumours </s> disease neuroendocrine tumours is caused by gene Cathepsin A causing disease Bone Marrow Failure </s> disease neuroendocrine tumours is caused by gene S2242 causing disease nasopharyngeal secretions </s>] You can deduce gene U2af1-rs1 <causing> disease neuroendocrine tumours.
YELP For phrase great taste: [phrase great taste indicate business Mika’s Greek rate stars 4.0 score business Thai Basil described with phrase shrimp risotto </s> phrase great taste indicate business Sonic Drive-In rate stars 5.0 score business Rubio’s described with phrase joe clark </s> phrase great taste indicate business Fired Pie described with phrase mind’s eye </s> phrase great taste indicate business Dj’s Bagel Cafe described with phrase free appetizer </s>] For business Sal’s Tuscan Grill: [business Sal’s Tuscan Grill rate stars 4.0 score business Bosa Donuts </s> business Sal’s Tuscan Grill rate stars 4.0 score business uBreakiFix </s> business Sal’s Tuscan Grill described with phrase iced americano indicate business El Palacio </s> business Sal’s Tuscan Grill described with phrase lamb vindaloo indicate business Chapman Volkswagen Scottsdale </s>] You can deduce phrase great taste <indicate> business Sal’s Tuscan Grill.
Appendix J Examples and Parsing Results
Examples and parsing results of GML and GraphML formats are provided in Figure 5.
Figure 5: Examples and parsing results of GML and GraphML formats.